What annual rate with continuous compounding would be required for a debt of $1,632 to grow into $3,348 in 19.5 years?

In this question you will need to solve for r in FV = PVert. Start by dividing both sides by PV. Then use logarithms to "bring down" the exponent.
Answer #1

PV =1632
FV =3348
time =19.5
FV =PV*e^rt
3348 =1632*e^(r*19.5)
3348/1632 =e^(r*19.5)
applying ln on both sides we get
ln(3348/1632)=r*19.5
r =ln(3348/1632)/19.5 =3.68%